# An End-to-End Smart Predict-then-Optimize Framework for Vehicle Relocation Problems in Large-Scale Vehicle Crowd Sensing

### v1

@@ Abstract
Ubiquitous mobile devices have catalyzed the development of vehicle crowd sensing (VCS). In particular, vehicle sensing systems show great potential in the flexible acquisition of extensive spatio-temporal urban data through built-in smart sensors under diverse sensing scenarios. However, vehicle systems like taxis often exhibit biased coverage due to the heterogeneous nature of trip requests and varying routes. To achieve a high sensing coverage, a critical challenge lies in how to optimally relocate vehicles to minimize the divergence between the spatio-temporal distributions of vehicles and target sensing distributions. Conventional approaches typically employ a two-stage predict-then-optimize (PTO) process: first predicting real-time vehicle distributions and subsequently generating an optimal relocation strategy based on the predictions. However, this approach can lead to suboptimal decision-making due to the propagation of errors from upstream prediction. To this end, we develop an end-to-end Smart Predict-then-Optimize (SPO) framework by integrating optimization into prediction within the deep learning architecture, and the entire framework is trained by minimizing the task-specific matching divergence rather than the upstream prediction error. Methodologically, we formulate the vehicle relocation problem by quadratic programming (QP) and incorporate a novel unrolling approach based on the Alternating Direction Method of Multipliers (ADMM) within the SPO framework to compute gradients of the QP layer, facilitating backpropagation and gradient-based optimization for end-to-end learning. The effectiveness of the proposed framework is validated using real-world taxi datasets in Hong Kong. Utilizing the alternating differentiation method, the general SPO framework presents a novel concept of addressing decision-making problems with uncertainty, demonstrating significant potential for advancing applications in intelligent transportation systems.
@@
journal: Transportation Research Part C: Emerging Technologies
@@ 1 Introduction
The MCS can be categorized into community sensing (CS) and vehicle crowd sensing (VCS) based on different types of sensor hosts (Ji et al. 2023a). CS utilizes sensors installed in mobile devices, while VCS leverages the sensing capabilities of vehicles to collect city-wide data (Xu et al. 2019). Host vehicles of VCS include taxis (Honicky et al. 2008), trams (Saukh et al. 2014), buses (Kang et al. 2016), and unmanned aerial vehicles (UAVs) (Li et al. 2022a). In general, the advantages of VCS in terms of temporal duration, geographical coverage, reliability, and consistency make it particularly suitable for scenarios requiring continuous, full coverage, and high-quality data (Ji et al. 2023a). Consequently, VCS is instrumental in domains that require high-precision data, such as infrastructure monitoring (Song et al. 2020, Hasenfratz et al. 2015, Liu and Ma 2022), traffic state estimation (Guo et al. 2022, Zhu et al. 2014) and traffic management (Ji et al. 2023b, Dai and Han 2023).
Specifically, the vehicles in the VCS framework are classified into two types: dedicated vehicles (DVs) and non-dedicated vehicles (NDVs). DVs, such as vehicles equipped with air pollution sensors and delivery drones with cameras, are fully guided by the dispatching center to prioritize sensing tasks. In contrast, NDVs, such as e-hailing vehicles and taxis, are under the complete control of their drivers. Upon receiving a sensing task, DVs will proceed immediately to the specific sensing locations. In contrast, NDVs operate under the control of drivers and are confined to their preferred cruising areas and routes. Moreover, NDVs can be converted into DVs through incentives such as monetary compensation.
@@
    An unrolling approach using alternating differentiation methods based on the ADMM for constrained optimization problems in large-scale networks. A novel unrolling approach is proposed to obtain the gradients of the constrained QP iteratively for large-scale networks. Specifically, the forward pass is solved using the ADMM, and the backpropagation is achieved through alternating differentiation.
@@
    Real-world experiments. Numerical experiments on a large-scale network with real-world data are conducted. Superior performance and unique characteristics of the proposed SPO framework are demonstrated.
The remainder of the paper is organized as follows: Section 2 reviews the related work on the vehicle relocation problem in mobile sensing. Section 3 introduces the SPO framework in vehicle relocation, including the prediction module, the optimization module, the integrated model, and an unrolling approach for the SPO framework. Section 4 presents the overall solution algorithm of the end-to-end SPO framework. Section 5 evaluates the proposed framework in real-world scenarios and compares the method with the baseline approaches. Section 6 concludes the results and proposes future directions.
@@ 2.1 Vehicle relocation in mobile sensing
Pervasive crowd sensing plays an important role in intelligent transport systems and smart city applications (Ji et al. 2023a, Dai and Han 2023). Given their relatively lower installation costs and extended operational duration, taxis are the most commonly utilized hosts in the VCS. However, due to heterogeneous trip patterns, the spatio-temporal distributions of these vehicles are often biased. To enhance sensing quality, numerous scholars have investigated intervention measures to effectively relocate vehicles through the development of incentivizing schemes (Fan et al. 2019, Xu et al. 2019, Chen et al. 2020) or proactive scheduling (Masutani 2015).

### v3

# SPO-VCS: An End-to-End Smart Predict-then-Optimize Framework with Alternating Differentiation Method for Relocation Problems in Large-Scale Vehicle Crowd Sensing

@@ Abstract
Ubiquitous mobile devices have catalyzed the development of vehicle crowd sensing (VCS). In particular, vehicle sensing systems show great potential in the flexible acquisition of extensive spatio-temporal urban data through built-in smart sensors under diverse sensing scenarios. However, vehicle systems like taxis often exhibit biased coverage due to the heterogeneous nature of trip requests and varying routes. To achieve a high sensing coverage, a critical challenge lies in how to optimally relocate vehicles to minimize the divergence between the spatio-temporal distributions of vehicles and target sensing distributions. Conventional approaches typically employ a two-stage predict-then-optimize (PTO) process: first predicting real-time vehicle distributions and subsequently generating an optimal relocation strategy based on the predictions. However, this approach can lead to suboptimal decision-making due to the propagation of errors from upstream prediction. To this end, we develop an end-to-end Smart Predict-then-Optimize (SPO) framework by integrating optimization into prediction within the deep learning architecture, and the entire framework is trained by minimizing the task-specific matching divergence rather than the upstream prediction error. Methodologically, we formulate the vehicle relocation problem by quadratic programming (QP) and incorporate a novel unrolling approach based on the Alternating Direction Method of Multipliers (ADMM) within the SPO framework to compute gradients of the QP layer, facilitating backpropagation and gradient-based optimization for end-to-end learning. The effectiveness of the proposed framework is validated using two real-world taxi datasets ranging from mid-size to large-scale in Hong Kong and Chengdu, China. Utilizing the alternating differentiation method, the general SPO framework presents a novel concept of addressing decision-making problems with uncertainty, demonstrating significant potential for advancing applications in logistics and intelligent transportation systems.
@@
journal: Transportation Research Part E: Logistics and Transportation Review
@@ 1 Introduction
The MCS can be categorized into community sensing (CS) and vehicle crowd sensing (VCS) based on different types of sensor hosts (Ji et al. 2023a). CS utilizes sensors installed in mobile devices, while VCS leverages the sensing capabilities of vehicles to collect city-wide data (Xu et al. 2019). Host vehicles of VCS include taxis (Honicky et al. 2008), trams (Saukh et al. 2014), buses (Kang et al. 2016), and unmanned aerial vehicles (UAVs) (Li et al. 2022a). In general, the advantages of VCS in terms of temporal duration, geographical coverage, reliability, and consistency make it particularly suitable for scenarios requiring continuous, full coverage, and high-quality data (Ji et al. 2023a).
Consequently, VCS is instrumental in domains that require high-precision data, such as infrastructure monitoring (Song et al. 2020, Hasenfratz et al. 2015, Liu and Ma 2022), traffic state estimation (Guo et al. 2022, Zhu et al. 2014) and traffic management (Ji et al. 2023b, Dai and Han 2023).
Specifically, the vehicles in the VCS framework are classified into two types: dedicated vehicles (DVs) and non-dedicated vehicles (NDVs). DVs, such as vehicles equipped with air pollution sensors and delivery drones with cameras, are fully guided by the dispatching center to prioritize sensing tasks. In contrast, NDVs, such as e-hailing vehicles and taxis, are under the complete control of their drivers and do not follow the guidance of the dispatching center. Upon receiving a sensing task, DVs will proceed immediately to the specific sensing locations. In contrast, NDVs operate under the control of drivers and are confined to their preferred cruising areas and routes. Moreover, NDVs can be converted into DVs through incentives such as monetary compensation.
@@
    An alternating differentiation approach based on the ADMM for constrained optimization problems in large-scale networks. A novel unrolling approach is proposed to obtain the gradients of the constrained QP iteratively for large-scale networks. Specifically, the forward pass is solved using the ADMM, and the backpropagation is achieved through alternating differentiation.
@@
    Two large-scale real-world experiments. Numerical experiments on two real-world datasets across five network sizes from mid-size to large-scale are conducted. Superior matching performance and computational efficiency of the proposed SPO framework are demonstrated.
The remainder of the paper is organized as follows: Section 2 reviews the related work on the vehicle relocation problem in mobile sensing. Section 3 introduces the SPO framework in vehicle relocation, including the prediction module, the optimization module, the integrated model, and an unrolling approach based on the alternating differentiation method for the SPO framework. Section 4 presents the overall solution algorithm of the end-to-end SPO framework. Section 5 evaluates the proposed framework in two real-world scenarios and compares the method with the baseline approaches. Section 6 concludes the results and proposes future directions.
@@ 2.1 Vehicle relocation in mobile sensing
Pervasive crowd sensing plays an important role in intelligent transport systems, logistics and smart city applications (Ji et al. 2023a, Dai and Han 2023, Yang et al. 2025a). Given their relatively lower installation costs and extended operational duration, taxis are the most commonly utilized hosts in the VCS. However, due to heterogeneous trip patterns, the spatio-temporal distributions of these vehicles are often biased. To enhance sensing quality, numerous scholars have investigated intervention measures to effectively relocate vehicles through the development of incentivizing schemes (Fan et al. 2019, Xu et al. 2019, Chen et al. 2020) or proactive scheduling (Masutani 2015).